OpenStudy (anonymous):

Solve x + 2y = 50 -2x + 3y = 40 5x + 5y + 5z = 300

OpenStudy (anonymous):

Can use elimination on the first two equations by multiplying the first equation by 2 and adding it to the second equation; that will let you solve for y. Put that value of y into the first equation to get x. With values for x and y, put both into the third equation to get z.

OpenStudy (anonymous):

Also, the third equation may be divided by 5 to reduce.

mathslover (mathslover):

x + 2y = 50 x = 50 -2y now put this value in second equation : \[\large{-2(50-2y)+3y=40}\] \[\large{-2(50)-2(-2y) + 3y = 40}\] \[\large{-100+4y+3y=40}\] \[\large{7y = 140}\] \[\large{y = \frac{140}{7}=20}\]

mathslover (mathslover):

now as we derived that ... x = 50-2y put y =20 x = 50 - 2(20) = 50 - 40 = 10 now put this in third equation : \[\large{5x+5y+5z=300}\] \[\large{5(10)+5(20)+5z=300}\] \[\large{50+100+5z=300}\] \[\large{5z=300-150}\]

mathslover (mathslover):

what do you get for z ?

mathslover (mathslover):

@DFlame28 please reply... what is z if 5z = 150..